# The statistical spread of transmission outages on a fast protection time scale based on utility data

### v1

@@ Abstract
When there is a fault, the protection system automatically removes one or more transmission lines on a fast time scale of less than one minute. The outaged lines form a pattern in the transmission network. We extract these patterns from utility outage data, determine some key statistics of these patterns, and then show how to generate new patterns consistent with these statistics. The generated patterns provide a new and easily feasible way to model the overall effect of the protection system at the scale of a large transmission system. This new generative modeling of protection is expected to contribute to simulations of disturbances in large grids so that they can better quantify the risk of blackouts. Analysis of the pattern sizes suggests an index that describes how much outages spread in the transmission network at the fast timescale.
@@ I Introduction
We present observed utility outage data that shows patterns in line outages at the time scale of less than one minute, at which the protection system acts. The line outages occurring during each minute mostly appear on the network as single lines or small tree subnetworks. The protection system routinely limits the outages to simple patterns, most frequently to outages involving one or two lines. However, much rarer but more impactful patterns with many line outages also appear in the data.
@@ II Literature review
There is extensive literature assessing the cascading risk of large transmission system blackouts by simulating models as surveyed in [8, 9]. Most of this literature outages lines without any detailed consideration of the protection system. However, there are advances in modeling some aspects of the protection system in this context.
@@
Rios et al. [10] model protection system misoperation by a constant probability when a fault lies in the vulnerability region of a relay. Chen et al. [11] model hidden protection failure of a line exposed by a neighboring outage with a probability that changes according to the line loading. Yu and Singh [12] obtain probabilities of failure to operate and the undesired trip from a steady state Markov chain and then use these probabilities to simulate cascades in a 24-bus power system. Dobson et al. [1] model the protection system in some detail and account for stuck breakers when cosimulating protection and power system dynamics in a 130-bus power system. Another application where detailed protection modeling has been applied is assessing resilience to the risk of extreme weather. Ciapessoni et al. [13] assess this risk in an 80-bus power system, and their models include busbar and double circuit faults and stuck breakers.
@@
Except for [4], all these studies advance the detailed modeling of specific mechanisms of protection operation and misoperation. These studies complement and have a different scope than our investigation of all the outage spreading that actually occurred at the fast time scale. Detailed modeling of specific mechanisms allows the effect of that mechanism to be assessed in simulations of blackout risk in small transmission system networks, whereas our approach based on observed data faithfully reflects the combined effect of all the mechanisms in large-scale transmission systems.
@@
Generative modeling of general cascading outages is used by Kelly-Gorham et al. [14, 15] as part of modeling transmission system resilience in the CRISP Computing Resilience of Infrastructure Simulation Platform. Starting from a random initial line outage, successive line outages are chosen according to the observed distribution of cascade size and the observed distribution of distances between successive line outages in resilience events. The line and generator restoration times are also sampled from their observed lognormal distributions. The generative modeling of line outages in CRISP has a similar overall objective to this paper: to generate outages consistent with the statistical patterns of observed events. However, it models all the cascading and outage processes combined and does not separately consider protection processes at the fast time scale. As this paper shows, fast protection processes have a special structure different from cascading or weather-induced outages in general. Most notably, outage propagation at the fast protection time scale produces small connected subnetworks that are mostly trees, whereas general cascading or weather-induced outages propagate locally, as well as more globally to disconnected sets of lines. This explains why this paper controls how the new lines are directly attached to the evolving patterns, whereas CRISP samples new lines according to their network distance from the initial line outage.
@@
While the only previous generative models we know in power transmission systems are in CRISP [14, 15], generative models of cascading phenomena have been used in other applications. Generative models have been used to model cascades of product recommendations in social networks and in discussion forums [16, 17]. For example, Leskovec et al. examine the topology of blog posts when some of them become popular and observe that the cascade size distribution presents a heavy tail and follows a Zipf distribution. They develop a generative model based on contagion dynamics that is able to accurately model the statistics of the observed cascades [16]. Gomez et al. introduce a more sophisticated model based on preferential attachment [17]. They derive a likelihood function used for parameter estimation and they fit their model to four different blogging communities. The model can replicate observed statistics, and also, the fitted parameters are used to draw conclusions about communication habits. A later review of generative models for online discussion threads points out the successes and diverse uses of these models, such as comparison of discussion platforms, predicting user behavior, and evaluation of platform design [3].
@@ III Utility data
For both transmission systems, the automatic line outage data include the outaged line specified by its sending and receiving end buses and the line outage times at a one-minute granularity. Other details of the outage data that are not used in this paper are described in [22] and [20]. Then, the automatic line outages are grouped together into the lines that outage within the time span of the same minute [^1] Repeats of outages of the same line in the same group are removed.. Each group of lines is called a “generation” of line outages in [22, 21, 4], although these references use a slightly different definition of generation [^2] References [22, 21, 4] define a generation as groups of outages that are separated by at least 1 minute, giving generations that sometimes include more outages. For example, if outages occur in each of 2 successive minutes, outages in each minute would be a generation in this paper, whereas all these outages would be included in one generation as defined in [22, 21, 4]. . Our outage time data has one-minute granularity, so the group is simply the outages with that minute recorded. Most of the groups of outages in one minute are a connected subnetwork of the transmission system network, and in that case the pattern is that group of outages. For a small fraction of cases (2.6% for BPA and 1.1% for NYISO), there is more than one connected component in the group. In that case, we consider each of the components in the group as a pattern, as combined failures stemming from fast protection actions can be assumed to result in a connected pattern.
@@ IV Observed patterns and their statistics
Patterns observed in the BPA and NYISO data are shown in Figs. 3 and 5, together with their bus degrees and a count of how many times those bus degrees appear in the data. (Recall that the bus degree is the number of lines incident at each bus and that the list of bus degrees is known as the degree sequence in network theory.) Most of the patterns are small trees, with only a few patterns with loops. For BPA, there are 11836 patterns (623 patterns per year) of which 93% have one line, 5% have two lines, and 2% have 3 or more lines. For NYISO, there are 7362 patterns (609 patterns per year) of which 93% have one line, 6% have two lines, and 1% have 3 or more lines.
The simpler patterns in Figs. 3 and 5 are much more common than the more complicated patterns, many of which only occur once or twice. This is expected in a well-engineered protection system, as it is designed to quickly halt the propagation of outages. The distributions of the number of lines in the patterns are shown by the red dots in the log-log plots in Fig. 2. These observed data in Fig. 2 can be fit with a Zipf distribution using the maximum likelihood method in [23]. The Zipf distribution (often called the zeta distribution or discrete Pareto distribution) is a heavy-tailed probability distribution $Z$ on the positive integers with
@@
as special cases. This gives an exact answer for the observed NYISO patterns. However, the situation is more complicated for more elaborate non-tree patterns in BPA because there are many possible orders in which the pattern can be assembled, and only the orders in which the line forming the loop is attached to join buses of degree 1 in the evolving pattern have double counting. Since there are only two more elaborate non-tree patterns for the observed patterns in BPA and none for NYISO, and there is an overcount by one in only some of the possible assembly orders, it is an underestimate that is a good approximation for the purpose of estimating $p_{1+}^{\rm observed}$ in (5) below to define $n_{1+}$ by (3) together with the special cases (4).
Note that we do not know from our data the initial line of each observed pattern or the order in which lines were added to obtain each observed pattern.
@@ V Generative model of fast protection patterns
The patterns are formed by starting with a single line out and successively adding lines until the process stops. The process is probabilistic so that it can be repeatedly sampled to produce a range of patterns that reproduce some key statistical features of the observed patterns. This process is only intended to reproduce these key statistical features of the observed patterns and is not intended to reflect the order of outages when the pattern is produced by the protection system. It aims to represent the statistics of the overall effect observed after the protection mechanisms have acted.
@@
Note that the objective is not to reproduce the observed patterns exactly but to be able to generate new patterns that are credible because they are statistically similar in key respects. If one wanted to use the exact patterns, then this could also be done directly from the specific historical patterns observed, but this lacks flexibility in the number and variety of samples, and since the larger patterns are rare, the historical patterns are only a limited sample of the possibilities. Our model is able to generate rare large patterns in new locations. There is a uniformity assumption that the same generative algorithm applies uniformly across the network, but this assumption is reasonable for this first generative model.
Two samples of 1000 generated patterns for each transmission system are shown in Figs. 4 and 6.
VI Protection Event Propagation System Index
@@
PEPSI is analogous to the System Event Propagation Index (SEPSI): PEPSI measures how many outages propagate at the fast protection time scale, whereas SEPSI measures how many generations of outages propagate in the network at a slower time scale. SEPSI is described in [24, 25].
@@ VIII-B Distance between distributions of degree sequences
Here, $\delta_{i}$ are the degree sequences of patterns, and $p_{i}$, and $q_{i}$ are the corresponding probabilities. The Wasserstein metric between these two discrete distributions computes the minimal transport plan that is required to convert one distribution into another based on the distance between degree sequences. For this particular case, the Wasserstein metric is the solution of a linear program
@@ VIII-C Distance between the degree sequences of observed and generated patterns
We evaluate the distance between the observed patterns and the same number of samples of the generated patterns by computing the Wasserstein distance between the distributions of their degree sequences. Because the generative model produces different sets of patterns every time it runs, there is variance in the evaluation of this distance. Thus, we perform this evaluation 100 times and compute statistics of the result.
For the BPA system, the mean distance between the generated and observed patterns is 0.0086 while its variance is 0.0016. The Wasserstein distance of 0.0086 corresponds to changing $0.0086\times 11836=102$ lines to change the 11836 generated pattern degree sequences into those observed. For the NYISO system, the mean distance of the generated and observed data is 0.0131 while its variance is 0.0027. The Wasserstein distance of 0.0131 corresponds to changing $0.0131\times 7362=96$ lines to change the 7362 generated pattern degree sequences into those observed.
@@ VIII-D Statistical testing of the observed and generated patterns
The generative model is used to produce 100 sets of patterns, with each set containing the same number of patterns as those observed (11836 patterns for BPA and 7362 patterns for NYISO). A permutation test is run on each of the 100 sets of patterns to test the null hypothesis that each set of patterns is from the same underlying distribution as the observed patterns. Each permutation test samples 10000 permutations. The p-value for each test is the estimated probability that the distance of the generated set of patterns from the observed set of patterns exceeds the calculated distance. We do not reject the hypothesis that the set of patterns is from the same underlying distribution if the p-value0.05absent0.05$\geq$ 0.05≥ 0.05.
This testing procedure gives 100 p-values for each transmission system. For BPA, the median p-value is 0.3565, and 98 p-values exceed 0.05, and for NYISO, the median p-value is 0.0983 and 73 p-values exceed 0.05. These results indicate that in most cases the samples from our generative model are indistinguishable from the observed data, thus strengthening the case that our model is able to capture the underlying statistical structure of the outage propagation.
@@ IX Discussion and Conclusions
We show how to extract network patterns in transmission line outages from utility outage data at the fast protection time scale and the large transmission system scale. The utility data is routinely collected by transmission operators and includes the bus names at each end of the line and the outage start times to the nearest minute. Given unique bus names, it is straightforward to form the network topology directly from the utility outage data itself so that the patterns of outages can be easily located on the network [21]. The outage patterns are easily extracted by noting which outages occur in the same minute.
@@
These key statistics can be sampled to generate representative sets of outage patterns consistent with the observed statistics. Indeed, this amounts to a novel generative model of the protection system’s overall effects at the transmission system scale.
@@
The ability to generate representative outage patterns from any given starting line outage, caused, for instance, by weather or by equipment failure, is important because the historical outages alone will not cover the full range of credible possibilities when assessing the risk of future outages. Since they are based on utility data, the generated patterns incorporate the common cases of routine operation as well as the rare, but more impactful complicated cases (protection backup, substation design, stuck breaker, hidden failure, common mode etc.) in which outages quickly spread further in the network. The patterns of outages on the network can be generated according to their statistics without getting involved in the various mechanisms for their cause and the formidable difficulties in practice of obtaining and representing the details of the protection system and substations across an entire transmission system.
@@
We show sample results of the generated patterns. To test these generated patterns quantitatively, we define distances between sets of patterns and show that the generated patterns are close to the observed patterns. Moreover, we consider whether the generated patterns are from a different probability distribution than the observed patterns when differences in their distance are evaluated. A statistical permutation test shows that it is unlikely that the generated patterns are from a different underlying probability distribution than the observed patterns.
@@
This paper is devoted to analyzing observed utility data and developing a new generative model of patterns of outages at the large transmission network scale and the fast protection system time scale. Promising future applications for the new generative model that further motivate its development include contributing a practical statistical model of the effects of protection that can be included in Monte Carlo simulations to help assess blackout risk at the transmission system scale, as well as providing observed data and statistics to validate or help guide the development of detailed protection models.

### v2

@@ Abstract
When there is a fault, the protection system automatically removes one or more transmission lines on a fast time scale of less than one minute. The outaged lines form a pattern in the transmission network. We extract these patterns from utility outage data, determine some key statistics of these patterns, and then show how to generate new patterns consistent with these statistics. The generated patterns provide a new and easily feasible way to model the overall effect of the protection system at the scale of a large transmission system. This new data-driven generative modeling of protection is expected to contribute to simulations of disturbances in large grids so that they can better quantify the risk of blackouts. Analysis of the pattern sizes suggests an index that describes how much outages spread in the transmission network at the fast timescale.
Index Terms:
Power transmission, Power system protection, resilience, reliability, statistics, generative AI
@@ I Introduction
We present observed utility outage data that shows patterns in line outages at the protection system time scale of less than one minute. The line outages occurring during each minute mostly appear on the network as single lines or small tree subnetworks. The protection system routinely limits the outages to simple patterns, most frequently to outages involving one or two lines. However, much rarer but more impactful patterns with many line outages also appear in the data.
@@ II Literature review
There is extensive literature assessing the cascading risk of large transmission system blackouts by simulating models as surveyed in [8, 9]. Most of this literature outages lines without any detailed consideration of the protection system. However, there are advances in modeling some aspects of the protection system in this context. Rios et al. [10] model protection system misoperation by a constant probability when a fault lies in the vulnerability region of a relay. Chen et al. [11] model hidden protection failure of a line exposed by a neighboring outage with a probability that changes according to the line loading. Yu and Singh [12] obtain probabilities of failure to operate and the undesired trip from a steady state Markov chain and then use these probabilities to simulate cascades in a 24-bus power system. Dobson et al. [1] model the protection system in some detail and account for stuck breakers when cosimulating protection and power system dynamics in a 130-bus power system. Another application where detailed protection modeling has been applied is assessing resilience to the risk of extreme weather. Ciapessoni et al. [13] assess this risk in an 80-bus power system, and their models include busbar and double circuit faults and stuck breakers.
Anders et al. [14] describes in detail the failure statistics of air blast breakers, including the distribution of time to failure for forced outages. Bollen [15] describes probabilities of a variety of breaker failures. Grant [16] extracts condition-dependent failure probabilities and rates from breaker data.
Protection actions with multiple outaged lines include common mode events. Mittelstadt et al. [17] find the mean time between common mode failures of BPA transmission lines and Keel et al. [18] describe the statistics of WECC transmission line common mode failures. In particular, Keel et al. [18] confirm that almost all the common mode events in WECC start in the same minute. Other data sources for common mode and dependent failures are reviewed in [19]. These common mode failure statistics can be expressed as failure rates and used in steady state reliability models [20, 21].
@@
Except for [4], all these studies advance the detailed modeling of specific mechanisms of protection operation and misoperation. These studies complement and have a different scope than our investigation of all the outage spreading that actually occurred at the fast time scale. Detailed modeling of specific mechanisms allows the effect of that mechanism to be assessed in simulations of blackout risk in small transmission system subnetworks, whereas our approach based on observed data faithfully reflects the combined effect of all the mechanisms in large-scale transmission systems.
@@
Generative modeling of general cascading outages is used by Kelly-Gorham et al. [22, 23] as part of modeling transmission system resilience in the CRISP Computing Resilience of Infrastructure Simulation Platform. Starting from a random initial line outage, successive line outages are chosen according to the observed distribution of resilience event sizes and the observed distribution of distances between successive line outages in resilience events. The line and generator restoration times are also sampled from their observed lognormal distributions. Similar generative modeling is used by Cheng et al. [24] to sample network cascading caused by an earthquake.
The generative modeling of line outages in CRISP and [24] has a similar overall objective to this paper: to generate outages consistent with the statistical patterns of observed events. However, it models all the cascading and outage processes combined and does not separately consider protection processes at the fast time scale. As this paper shows, fast protection processes have a special structure different from cascading or weather-induced outages in general. Most notably, outage propagation at the fast protection time scale produces small connected subnetworks that are mostly trees, whereas general cascading or weather-induced outages propagate locally, as well as more globally to disconnected sets of lines. This explains why this paper controls how the new lines are directly attached to the evolving patterns, whereas CRISP and [24] sample new lines according to their network distance from the initial line outage.
@@
While the only previous generative models we know of in power transmission systems are in [22, 23, 24], generative models of cascading phenomena have been used in other applications. Generative models have been used to model cascades of product recommendations in social networks and in discussion forums [25, 26]. For example, Leskovec et al. examine the topology of blog posts when some of them become popular and observe that the cascade size distribution presents a heavy tail and follows a Zipf distribution. They develop a generative model based on contagion dynamics that is able to accurately model the statistics of the observed cascades [25]. Gomez et al. introduce a more sophisticated model based on preferential attachment [26]. They derive a likelihood function used for parameter estimation and they fit their model to four different blogging communities. The model can replicate observed statistics, and the fitted parameters are used to draw conclusions about communication habits. A later review of generative models for online discussion threads points out the successes and diverse uses of these models, such as comparison of discussion platforms, predicting user behavior, and evaluation of platform design [3].
@@ III Utility data
For both transmission systems, the automatic line outage data include the outaged line specified by its sending and receiving end buses and the line outage times at a one-minute granularity. Other details of the outage data that are not used in this paper are described in [31] and [29]. Then, the automatic line outages are grouped together into the lines that start their outage within the time span of the same minute [^1] Repeats of outages of the same line in the same group are removed.. Each group of lines is called a “generation” of line outages in [31, 30, 4], although these references use a slightly different definition of generation [^2] References [31, 30, 4] define a generation as groups of outages that are separated by at least 1 minute, giving generations that sometimes include more outages. For example, if outages occur in each of 2 successive minutes, outages in each minute would be a generation in this paper, whereas all these outages would be included in one generation as defined in [31, 30, 4]. .
Our outage time data have one-minute granularity, so the group is simply the outages with that minute recorded as their start time. Most of the groups of outages in one minute are a connected subnetwork of the transmission system network, and in that case the pattern is that group of outages. For a small fraction of cases (2.6% for BPA and 1.1% for NYISO), there is more than one connected component in the group. In that case, we consider each of the components in the group as a pattern, as combined failures stemming from fast protection actions can be assumed to result in a connected pattern. In summary, the patterns are the connected automatic outages that start at the same minute. It is easy to automatically process the outages to group them into patterns.
There are a few external causes that can cause multiple nearby faults at the fast time scale of less than one minute. For example, one of the large patterns observed in the BPA data is caused by an earthquake. Lightning could locally cause multiple faults at the fast time scale, either by multiple ground contacts of a single cloud to ground strike [32], or multiple cloud to ground strikes [33].
Groups of patterns of various types can be further analyzed. For example, the top 6 most common dispatcher cause codes for the BPA outages in patterns with 3 or more outages are: Lightning 15%, Foreign Trouble 15%, RAS Initiated 12%, Unknown 11%, Forced (Configuration) 10%, Terminal Equipment Failure 6%. If the line physical location is known, allowing weather data to be related to the pattern, then weather causes can be analyzed more reliably than with cause codes.
@@ IV Observed patterns and their statistics
Patterns observed in the BPA and NYISO data are shown in Figs. 2 and 3, together with their bus degrees and a count of how many times those bus degrees appear in the data. (Recall that the bus degree is the number of lines incident at each bus, and that the list of bus degrees is known as the degree sequence in network theory.) Most of the patterns are small trees, with only a few patterns with loops. For BPA, there are 11836 patterns (623 patterns per year) of which 93% have one line, 5% have two lines, and 2% have 3 or more lines. For NYISO, there are 7362 patterns (609 patterns per year) of which 93% have one line, 6% have two lines, and 1% have 3 or more lines.
The simpler patterns in Figs. 2 and 3 are much more common than the more complicated patterns, many of which only occur once or twice. This is expected in a well-engineered protection system, as it is designed to quickly halt the propagation of outages. The distributions of the number of lines in the patterns are shown by the red dots in the log-log plots in Fig. 4. These observed data in Fig. 4 can be fit with a Zipf distribution using the maximum likelihood method in [34]. The Zipf distribution (often called the zeta distribution or discrete Pareto distribution) is a heavy-tailed probability distribution $Z$ on the positive integers with
@@
as special cases. This gives an exact answer for the observed NYISO patterns. However, the situation is more complicated for the more elaborate non-tree patterns in BPA because there are many possible orders in which the pattern can be assembled, and only the orders in which the line forming the loop is attached to join buses of degree 1 in the evolving pattern have double counting. Since there are only two more elaborate non-tree patterns for the observed patterns in BPA and none for NYISO, and there is an overcount by one in only some of the possible assembly orders, it is an underestimate that is a good approximation for the purpose of estimating $p_{1+}^{\rm observed}$ in (5) below to define $n_{1+}$ by (3) together with the special cases (4). Note that we do not know from our data the initial line of each observed pattern or the order in which lines were added to obtain each observed pattern.
@@ V Generative model of fast protection patterns
The patterns are formed by starting with a single line out and successively adding lines until the process stops. The process is probabilistic so that it can be repeatedly sampled to produce a range of patterns that reproduce some key statistical features of the observed patterns. This process is only intended to reproduce these key statistical features of the observed patterns and is not intended to reflect the order of outages when the pattern is produced by the protection system. It aims to represent the statistics of the overall effect observed after all the protection mechanisms have acted.
@@
Note that the objective is not to reproduce the observed patterns exactly but to be able to generate new patterns that are credible because they are statistically similar in key respects. If one wanted to use the exact patterns, then this could also be done directly from the specific historical patterns observed, but this lacks flexibility in the number and variety of samples, and since the larger patterns are rare, the historical patterns are only a limited sample of the possibilities. Our model is able to generate rare large patterns in new locations. There is a uniformity assumption that the same generative algorithm applies uniformly across the network, but this assumption is reasonable for this first generative model. Sometimes, it may be useful to consider simplified models that restrict the patterns generated. For example, the very simplest model adds a second adjacent line with probability 7% to any given single line outage.
Two samples of 1000 generated patterns for each transmission system are shown in Figs. 6 and 7.
Since the generative model randomly selects an initial line and only adds some adjacent lines probabilistically, its execution is very fast and largely independent of network size.
VI Protection Event Propagation Slope Index
@@
PEPSI is analogous to the System Event Propagation Index (SEPSI): PEPSI measures how much outages propagate at the fast protection time scale, whereas SEPSI measures how much generations of outages propagate in the network at a slower time scale. SEPSI is described in [35, 36].
Conventional protection system reliability metrics generally describe the probabilities or annual rates of specific types of misoperation or operation for specific components or subsystems [37]. For example, breaker failures are quantified in [15, 14, 16]. The PEPSI metric is different in that it measures the overall spread of protection actions on an entire network for all causes of additional lines outaging.
@@ VIII-B Distance between distributions of degree sequences
Here, $\delta_{i}$ is the degree sequences of pattern $i$, and $p_{i}$ or $q_{i}$ are the corresponding probabilities. The Wasserstein metric between these two discrete distributions computes the minimal transport plan that is required to convert one distribution into another based on the distance between degree sequences. For this particular case, the Wasserstein metric is the solution of a linear program
@@ VIII-C Distance between the degree sequences of observed and generated patterns
We evaluate the distance between the observed patterns and the same number of samples of the generated patterns by computing the Wasserstein distance between the distributions of their degree sequences. Because the generative model produces different sets of patterns every time it runs, there is variance in the evaluation of this distance. Thus, we perform this evaluation 1000 times and compute statistics of the result.
For the BPA system, the mean distance between the generated and observed patterns is 0.009 while its standard deviation is 0.002. The Wasserstein distance of 0.009 corresponds to changing $0.009\times 11836=107$ lines to change the 11836 generated pattern degree sequences into those observed. For the NYISO system, the mean distance of the generated and observed data is 0.013 while its standard deviation is 0.003. The Wasserstein distance of 0.013 corresponds to changing $0.013\times 7362=99$ lines to change the 7362 generated pattern degree sequences into those observed. We regard these distances for BPA and NYISO as satisfactorily small.
@@ VIII-D Statistical testing of the observed and generated patterns
The generative model is used to produce 1000 sets of patterns, with each set containing the same number of patterns as those observed (11836 patterns for BPA and 7362 patterns for NYISO). A permutation test is run on each of the 100 sets of patterns to test the null hypothesis that each set of patterns is from the same underlying distribution as the observed patterns. Each permutation test samples 10 000 permutations. The p-value for each test is the estimated probability that the distance of the generated set of patterns from the observed set of patterns exceeds the original unpermuted distance. We do not reject the hypothesis that the set of patterns is from the same underlying distribution if the p-value0.05absent0.05\,$\geq$ 0.05≥ 0.05.
This testing procedure gives 1000 p-values for each transmission system. For BPA, the median p-value is 0.31, 97% of the p-values exceed 0.05, and 99.9% of the p-values exceed 0.01, and for NYISO, the median p-value is 0.08, 69% of the p-values exceed 0.05, and 93% of the p-values exceed 0.01. These results indicate that in almost all cases for BPA and in most cases for NYISO, the samples from our generative model are indistinguishable from the observed data, supporting the case that our model can often capture the underlying statistical structure of outage propagation.
The comparison between the generated and observed patterns is necessarily statistical in nature. While the permutation test inherently can only fail to disprove (rather than prove) the hypothesis that the generated and observed patterns are drawn from the same statistical distribution, the permutation test results comparing the closeness of the degree sequences of the generated and observed patterns are encouraging. The degree sequence is a significant characteristic of a pattern and the distance between degree sequences incorporates the intuition that close patterns differ by changing only a few lines. However, there could potentially be other degrees of freedom not constrained by the distances between degree sequences that could show some significant differences between the generated and observed patterns.
The deviations from the observed patterns in the generative model include the random aleatory variations when sampling from the generative model as well as epistemic imperfections in the form of the generative model, which may include the limited number of statistical characteristics reproduced by the generative model and a uniformity assumption. By using the data for the limited number of larger historical patterns to generate statistically similar patterns across the network, the generative model relies on an assumption that there is sufficient uniformity across the network in the statistics of fast failure propagation. This is a reasonable assumption for a transmission system in one specific geographic region that is designed and maintained by a single regional entity.
@@ IX Discussion and Conclusions
We show how to extract network patterns in transmission line outages from utility outage data at the fast protection time scale and the large transmission system size scale. The utility data include the bus names at each end of the line and the outage start times to the nearest minute. The outage patterns are the connected groups of lines that outage starting in the same minute. It is easy to note which outages start in the same minute to form the patterns. Then the patterns can be located on the network. The required data is systematically collected by all transmission utilities across North America and reported to NERC in the TADS Transmission Availability Data System. Similar data is collected world wide. Utilities know their own network topology so that the outage patterns can be located on their network. Alternatively, as done in this paper, given unique bus names, it is straightforward to form the network topology directly from the utility outage data itself so that the patterns of outages can be located on the network [30]. We conclude that the network patterns can easily be found from data that is generally collected by transmission utilities worldwide.
@@
These key statistics can be sampled to generate representative sets of outage patterns consistent with the observed statistics. Indeed, this amounts to a novel generative model of the protection system’s overall effects at the transmission system scale. The generative model is inherently computationally fast, producing network patterns very quickly even in large networks.
@@
The ability to generate representative outage patterns from any given starting line outage, caused, for instance, by weather or by equipment failure, is important because the historical outages alone will not cover the full range of credible possibilities when assessing the risk of future outages. Since they are based on utility data, the generated patterns incorporate the common cases of routine operation as well as the rare, but more impactful complicated cases (protection backup, substation design, stuck breaker, hidden failure, common mode etc.) in which outages quickly spread further in the network. The patterns of outages on the network can be generated according to their statistics without getting involved in the various mechanisms for their cause and the formidable difficulties in practice of obtaining and representing the details of the protection system and substations across an entire transmission system. It is best to apply the generative model to a transmission system with the parameters obtained from the detailed outage data for that transmission system. However, if the detailed outage data is not available, as for example in IEEE test systems or synthetic grids, the generative model could be applied with parameters typical of those observed in other systems. The generative model is inherently computationally fast, producing network patterns very quickly even in large networks.
@@
We show sample results of the generated patterns. To test these generated patterns quantitatively, we define distances between sets of patterns and show that the generated patterns are close to the observed patterns. Moreover, we consider whether the generated patterns are from a different probability distribution than the observed patterns when differences in their distance are evaluated. A statistical permutation test shows that it is unlikely in most cases that the generated patterns are from a different underlying probability distribution than the observed patterns.
@@
This paper is devoted to analyzing observed utility data and developing a new generative model of patterns of outages at the large transmission network scale and the fast protection system time scale. The required data is generally available to transmission utilities and the required computations are straightforward, fast, and scalable, so that finding the observed patterns and generating statistically similar patterns is widely applicable. Promising future applications for the new generative model include contributing a practical statistical model of the effects of protection that can be included in Monte Carlo simulations assessing blackout risk at the transmission system scale, providing observed data to calibrate other protection models, and generating enough synthetic data for rare protection events to train AI models.
ACKNOWLEDGEMENT
We gratefully thank Bonneville Power Administration and New York Independent System Operator for making publicly available the outage data that made this paper possible. The analysis and any conclusions are strictly those of the authors and not of BPA or NYISO.
Government License (will be removed at publication): The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan.